### v1

@@ Abstract
Large language models (LLMs) are revolutionizing every aspect of society. They are increasingly used in problem-solving tasks to substitute human assessment and reasoning. LLMs are trained on what humans write and thus prone to learn human biases. One of the most widespread human biases is overconfidence. We examine whether LLMs inherit this bias. We automatically construct reasoning problems with known ground truths, and prompt LLMs to assess the confidence in their answers, closely following similar protocols in human experiments. We find that all five LLMs we study are overconfident: they overestimate the probability that their answer is correct between 20% and 60%. Humans have accuracy similar to the more advanced LLMs, but far lower overconfidence. Although humans and LLMs are similarly biased in questions which they are certain they answered correctly, a key difference emerges between them: LLM bias increases sharply relative to humans if they become less sure that their answers are correct. We also show that LLM input has ambiguous effects on human decision making: LLM input leads to an increase in the accuracy, but it more than doubles the extent of overconfidence in the answers.
@@
Large language models (LLMs) are revolutionizing every aspect of society. Among other things, they are used in problem-solving tasks that require careful reasoning and assessment. While LLMs are well-known to reliably reproduce knowledge on which they have been trained (Zhao et al., 2023), they are prone to making mistakes in reasoning tasks on which they have not been directly trained (Li et al., 2024b; Huang and Chang, 2023; Chang et al., 2024; Xu et al., 2023; Pawitan and Holmes, 2025).
@@
LLMs are trained on what humans write. As such, they are prone to develop human biases in their reasoning. One of the most prevalent biases in human judgment is overconfidence (Hoffrage, 2022; Kahneman, 2011; Moore and Healy, 2008; Malmendier and Taylor, 2015). Humans are prone to underestimating the limits of their knowledge (Dunning, 2011; Kruger and Dunning, 1999), leading them to overestimate the accuracy of their knowledge. Evaluating the correctness of their reasoning in different domains may also lead to the conjunction fallacy (Tversky and Kahneman, 1983), amplifying the overall overconfidence in judgment.  [^1] Various non-cognitive factors have also been shown to contribute to overconfidence, such as social signaling, (Burks et al., 2013), environmental factors (Goette et al., 2015), or strategic considerations to feign confidence (Charness et al., 2018) that may also provide an evolutionary mechanisms for the trait (Johnson and Fowler, 2011; Marshall et al., 2013).
@@
Overconfidence has been shown to affect decision in many areas. In the professional domain, overconfident managers are more likely to use equity financing (Malmendier and Tate, 2005; Malmendier et al., 2011) and more likely to engage in unprofitable merger (Malmendier and Tate, 2008), exhibiting little awareness or learning over time (Huffman et al., 2022). There is also evidence of overconfidence affecting everyday behaviors, such as working out regularly or dietary choices (Arni et al., 2021), or financial investments (Barber and Odean, 2001).
@@
In this context, LLMs may hold, both, promise and peril. On the one hand, overconfident individuals may be challenged in their reasoning by LLM input, thus mitigating human bias. On the other hand, if overconfidence manages to enter LLMs, this may exacerbate confidence in human decision makers.
@@
In this paper, we examine whether LLMs inherit overconfidence when evaluating their own reasoning, and whether this affects human decision makers. While conventional wisdom starts from the "bias in, bias out" hypothesis (Barocas and Selbst, 2016; Mayson, 2019), this need not always hold: theoreitcal work shows that overconfident agents learning from mis-specified models have complex and sometimes counterintutive implications heidhues2019overconfidence. It is possible that biased training data leads AI to be less biased towards that group (rambachan2020bias). 222rambachan2020bias show that in many realistic environments, bias in training data, stemming, e.g. from human bias against demographic group, can lead to less biased behavior of an AI towards that group. However, a corollary of their result is that biased training data will still affect other groups, with the AI being more biased towards them, even though humans don’t display bias towards those groups. To our knowledge, no formal results exist for the specific context of AI or LLM reasoning when trained on biased data. We thus test for overconfidence in LLM reasoning, and probe into the underlying mechanisms. We ask whether the bias is amplified or weakened relative to humans performing the same task. We prompt five commonly used LLMs with a large number of reasoning problems to which we know the true answer. The questions are generated using the algorithm in li2024drowzee to generate questions on which the LLMs are unlikely to be trained on. Thus, to arrive at an answer, the LLMs need to engage in reasoning on the spot.  [^3] Several recent papers have examined measures of calibration that can be extracted from LLMs’ intermediate outputs. Prior work has taken three main approaches to measuring LLM confidence: (1) logit-based estimation, which requires internal model access (yin2023large; gpt4report2023; he2023investigating; zhang2024calibrating); (2) direct confidence elicitation through prompting (tian2023just; xiong2024can; wei2024measuring; wen2024mitigating); and (3) auxiliary model approaches, ranging from single-model prediction (ulmer2024calibrating)to multi-source integration methods (zhao2024pareto). These calibration studies predominantly rely on standard question-answering datasets, and it is likely that the LLMs have been trained on a large fraction of the questions used (tian2023just; wei2024measuring; zhang2024calibrating; xiong2024can). Thus, these studies do not address LLMs’ confidence in questions that require them to reason. By contrast, we utilize algorithmically generated questions that guarantee minimal training contamination, allowing us to study confidence in answers that involve reasoning by LLMs. We ask the LLMs to assess the confidence in their answers, closely following similar protocols in human experiments. We elicit the LLM’s confidence in the correctness of the answer, the facts used, and the reasoning.
@@
We find that all five LLMs we examine are overconfident (Table 1). The LLMs on average overestimate the probability that their answer is correct between 20% (for GPT o1) and 60% (for GPT 3.5), with the other models falling somewhere in between. We find that more advanced models such as GPT 4o or GPT o1 have higher accuracy than GPT 3.5 or Llama 3.2, in line with earlier findings (shahriar2024putting; zhong2024evaluation; dubey2024llama). Interestingly, the confidence judgments across models are very similar, and in no way reflect the large differences in accuracy rates.
@@
Next, we examine how, within each LLM model, accuracy is related to confidence across the different questions. Panels A and B of Figure 1 show the accuracy rates as a function of confidence, for the GPT and Llama models, respectively. The panel shows a strong and positive association between accuracy and confidence. The graphs also show that even when a model is fully confident in its answer, there is still substantial bias: Panel A of Table 2 shows that the bias varies between 14% and 20% for the top performing models (GPT 4o, o1, and Llama 3.1), and even more for GPT 3.5 and Llama 3.2.
@@
For all models except Llama 3.2, bias increases with lower confidence levels. The estimates in Table 2 quantify this effect: a 10% drop in confidence is associated with a drop in accuracy of around 25% for GPT 4o and GPT o1, and around 15% for Llama 3.1 (see panel A in Table 2). This implies that the bias gets larger as the models’ confidence declines.
@@
The strong association between confidence and accuracy also extends each component of the reasoning process in which the LLMs have to engage. For each question, we separately elicit the the LLM’s confidence that the facts used are correct, and that the reasoning applied is correct. These confidence measures are highly correlated with confidence that the overall answer is correct (correlations range from 0.49 to 0.92 across models , see Table D.7 and Table D.7 in the SI). We also calculate the semantic similarity between the facts used and the reasoning applied with the sufficient set of facts, and confidence that the reasoning is correct (see section A.3 in the SI for details). We find significant associations between confidence and similarity for, both, facts and reasoning ($p<0.001$ in all cases, see Tables D.7 and D.7 in the SI). Thus, there is a palpable, but biased, relationship between accuracy and confidence in the answers, and this association extends to the necessary components for the reasoning.
@@
We also examined the stability of these results. In five replications of the results for GPT 3.5 and 4o, we retrieve nearly identical results with regard to the answer and the confidence. The answers change in less than 2 percent in each iteration. The changes we observe also do not have a clear direction: overall, accuracy only increase by 1.7 percent after five replications (see section D.8 in the SI).
@@
We compare the accuracy, confidence, and Dunning-Kruger effect of LLMs to a human benchmark. In a large-scale online experiment on Prolific, we recruit participants to answer 10 questions of a subset of 2000 (see SI section B). Like the LLMs, the human participants have to pick one of the answers (Yes or No), and indicate the probability with which they believe that their answer is correct. Human participants answer 66% of the questions correctly. This puts them on par in terms of accuracy with GPT 4o or Llama 3.1, but slightly behind GPT o1. Their confidence is significantly lower than that of any of the LLMs: human participants think their accuracy is 70%, thus overestimating it by only 4% on average ($p<0.01$, see Table S2 in the SI). This is far lower overconfidence than any of the LLMs with comparable accuracy. There is a maked difference in shape of the distribution of confidence (see Figure S5 in the SI): compared to LLMs, humans’ confidence appears far more nuanced, and with a mode (nearly 35% of the answers) stating that it is 50:50 whether their answers are correct.
@@
Panel C of Figure 1 shows the relationship between accuracy and confidence in the human sample, compared to the three most accurate LLMs. Two striking features emerge from the graph: first, when humans are 100% sure that their answer is correct, their accuracy rate is significantly lower at 81% ($p<0.01$, see column 1 in Table S11 ). Thus, for questions that humans consider relatively easy and are thus 100% sure of their answer, they are highly overconfident. This level of overconfidence is roughly on par with the same conditional accuracy rates of GPT 4o or Llama 3.1. However, while originating from the same level of bias when 100% sure of their answers, the confidence gradients between LLMs and humans are sharply different. As discussed before, LLMs become more biased as they enter more uncertain territory. In sharp contrast, the human confidence gradient is only about 0.5, indicating that human bias decreases as humans become less certain. As can be seen in the figure, humans eventually end up slightly underconfident: when participants think their accuracy is only 50%, it is actually slightly higher (54%, significantly higher than they expect $p<0.01$).
@@
We interpret this sharp difference in the confidence gradient as a stark manifestation of the Dunning-Kruger effect (dunning2011dunning; kruger1999unskilled). Our test cases require LLMs to reason to answer them, as they are unlikely to be trained on the answer. This may cause LLMs being "trapped" in their prediction model in a way that is hard for them to overcome: the model-internal mechanisms cannot provide them with a sense of what knowledge is not part of their training data. The LLMs pick the answer that is most consistent with what they were trained on, and form an estimate of their accuracy based on what is in their training data. As LLMs move into less certain territory, this problem gets stronger, and causes LLMs to have much less of a sense of the limits of their knowledge. Human participants may recognize in a question, e.g., that one of the names is a Roman emperor, and realize that she does not know much about Roman history. By contrast, in an LLM, there is no mechanism to tell it that there might be other Roman public figures that it wasn’t trained on. This results in overconfidence far stronger than what is observed in humans though the mechanisms of the Dunning-Kruger effect. Interestingly, while more advanced models like GPT 4o and GPT o1 have higher accuracy rates for test cases where they are certain their answer is correct, they also display a significantly stronger confidence gradient, and, hence a stronger Dunning-Kruger effect.
@@ Discussion
The implications of LLM overconfidence span practical, ethical and technical dimensions across stakeholder groups. Our results provide users with a new benchmark for how users of LLMs should approach reasoning from LLMs, counteracting the widespread overestimation of LLM capabilities (KLINGBEIL2024108352; holbrook2024; info:doi/10.2196/56764). Secondly, our results provide guidance for the future development of LLMs. Architectural choices in training objectives currently prioritize fluency over accuracy, incentivizing fabricated but coherent responses (yin2023large; fi15100336). The lack of a built-in uncertainty correction mechanism generates hallucinated responses and overconfidence (shorinwa2024survey). Current evaluation metrics (wei2024measuring; geng-etal-2024-survey) fail to capture calibration quality, necessitating new frameworks for quantifying and gauging overconfidence such as the one we presented here. Developing novel interpretability frameworks (wen2024from; 9857619) that pinpoint how and why overconfident and biased predictions emerge can support clearer communication of model uncertainty and limitations to users.
@@
Model a) Fraction of b) LLM confidence c) Bias calculated
correct answers in answer ($\widetilde{p}$) based on $\widetilde{p}$
@@
  • Notes: $N=10,000$. Standard deviations in parentheses (divide by 100 to obtain approximate SE of means). All accuracy rates and confidence measures are significantly different between models. All biases are significantly different from zero. ($p<0.001$, see SI Table S9.)
@@
Predicted accuracy 0.43 0.80 0.86 0.83 0.65
if $\widetilde{p}=1$ (0.01) (0.01) (0.01) (0.01) (0.01)
$N$ 9,795 9,973 9,990 9,750 9,910
$R^{2}$ 0.01 0.07 0.07 0.08 0.01
@@
  • Notes: Dependent variable is correct answer (= 1). The table presents the gradient of accuracy with respect to self-reported confidence in answer correctness across models. Confidence levels are normalized by subtracting 1 from the original confidence scores. Regression is weighted by the frequency of confidence levels. All coefficients are significant at p<0.01.
Human Baseline Confidence LLM Confidence
$\Delta$ Accuracy $\Delta$ Bias $\Delta$ Accuracy $\Delta$ Bias
@@
  • Notes: LLM confidence is mean-centered at 0.8; Baseline human confidence is mean-centered at 0.5; High Confidence is defined as confidence above 0.9 for LLM and above the median for human baseline. The regressions also control for high confidence indicators, though these results are not displayed in the table. Standard errors in parentheses are clustered at the participant and question levels. Illustrations can be seen in SI Figure S4.
@@
Notes: Panel A shows the relationship between accuracy and confidence for GPT models, and Panel B shows the same relationship for Llama models. Panel C compares the accuracy-confidence relationship between advanced LLM models and human responses on the question subset for the human baseline benchmark. The size of circles in each panel scales with the number of observations, and all slopes are significant ($p<0.001$ in all cases; see Table 2 for Panel A and B, and Table S11 in Supplementary Information for Panel C). Shaded areas indicate 95% confidence bands.
@@
Notes: Panel A shows the treatment effects on change in accuracy across experimental conditions. Panel B shows the treatment effects on the change in bias across the same conditions. Error bars represent 95% confidence intervals. For detailed regression analysis, see SI Table S12.

### v2

@@ Abstract
Large language models (LLMs) are revolutionizing every aspect of society. They are increasingly used in problem-solving tasks to substitute human assessment and reasoning. LLMs are trained on what humans write and are thus exposed to human bias. We evaluate whether LLMs inherit one of the most widespread human biases: overconfidence. We algorithmically construct reasoning problems with known ground truths. We prompt LLMs to answer these problems and assess the confidence in their answers, closely following similar protocols in human experiments. We find that all five LLMs we study are overconfident: they overestimate the probability that their answer is correct between 20% and 60%. Humans have accuracy similar to the more advanced LLMs, but far lower overconfidence. Although humans and LLMs are similarly biased in questions which they are certain they answered correctly, a key difference emerges between them: LLM bias increases sharply relative to humans if they become less sure that their answers are correct. We also show that LLM input has ambiguous effects on human decision making: LLM input leads to an increase in the accuracy, but it more than doubles the extent of overconfidence in the answers.
@@
Large language models (LLMs) are revolutionizing every aspect of society. Among other things, they are used in problem-solving tasks that require careful reasoning and assessment. While LLMs are well-known to reliably reproduce knowledge on which they have been trained (Zhao et al., 2023), they are prone to making mistakes in reasoning tasks on which they have not been directly trained (Li et al., 2024b; Huang and Chang, 2023; Chang et al., 2024; Xu et al., 2023; Pawitan and Holmes, 2025) and their reasoning may collapse completely if problems become too complex (Shojaee et al., 2025).
@@
LLMs are trained on what humans write. This exposes them to human bias, and puts them at risk to develop them in their reasoning. One of the most prevalent biases in human judgment is overconfidence (Hoffrage, 2022; Kahneman, 2011; Moore and Healy, 2008; Malmendier and Taylor, 2015). Humans are prone to underestimating the limits of their knowledge (Dunning, 2011; Kruger and Dunning, 1999), leading them to overestimate the quality of their judgments.  [^1] Various non-cognitive factors have also been shown to contribute to overconfidence, such as social signaling, (Burks et al., 2013), environmental factors (Goette et al., 2015), or strategic considerations to feign confidence (Charness et al., 2018) that may also provide an evolutionary mechanisms for the trait (Johnson and Fowler, 2011; Marshall et al., 2013).
@@
Overconfidence has been shown to affect decision making in many areas. In the professional domain, overconfident managers are more likely to use equity financing (Malmendier and Tate, 2005; Malmendier, Ulrike et al., 2011) and more likely to engage in unprofitable mergers (Malmendier, Ulrike and Tate, Geoffrey, 2008), exhibiting little awareness or learning over time (Huffman et al., 2022). There is also evidence of overconfidence affecting everyday behaviors, such as working out regularly or dietary choices (Arni et al., 2021), or financial investments (Barber and Odean, 2001).
@@
In this context, LLMs may hold, either, promise or peril. On the one hand, overconfident individuals may be challenged in their reasoning by LLM input, thus mitigating human bias. On the other hand, if overconfidence manages to enter LLMs, this may exacerbate bias in human decision makers.
@@
We first examine whether LLMs exhibit overconfidence in their own reasoning abilities. There are no general theoretical results to characterize in which contexts AI or LLM reasoning leads to biased answers. Conventional wisdom suggests the "bias in, bias out" hypothesis (Barocas and Selbst, 2016; Mayson, 2019): if models are trained on biased data, it is hypothesized that this biases its own responses in a similar way. However, this need not always hold. It is possible that training data exhibiting bias with regard to one attribute leads AI to be less biased towards the same attribute(Rambachan and Roth, 2020)  [^2] Rambachan and Roth (2020) show that in many realistic environments, bias in training data, stemming, e.g. from human bias against demographic group, can lead to less biased behavior of an AI trained on that data towards the group that was discriminated against in the training data. However, a corollary of their result is that biased training data will still affect other groups, with the AI being more biased towards them, even though humans don’t display bias towards those groups. Inference and reasoning can further be affected by overconfidence: e.g., Heidhues et al. (2018) show that overconfident, but otherwise rational, agents may develop persistent bias instead of learning from feedback. We thus aim to design an experiment to detect departures from unbiased calibration, and probe into the underlying mechanisms. We also designed our experiments such that they are sufficiently powered under the significance levels proposed by Benjamin et al. (2018) for novel findings.
@@
To this end, we prompt five commonly used LLMs with 10,000 test cases to which we know the true answer. The questions are generated using the algorithm in Li et al. (2024b) to generate questions on which the LLMs are unlikely to be trained on. Thus, to arrive at an answer, the LLMs need to engage in reasoning on the spot.  [^3] Several recent papers have examined measures of calibration that can be extracted from LLMs’ intermediate outputs. Prior work has taken three main approaches to measuring LLM confidence: (1) logit-based estimation, which requires internal model access (Yin et al., 2023; OpenAI, 2023; He et al., 2023; Zhang et al., 2024); (2) direct confidence elicitation through prompting (Tian et al., 2023; Xiong et al., 2024; Wei et al., 2024; Wen, Bingbing et al., 2024); and (3) auxiliary model approaches, ranging from single-model prediction (Ulmer et al., 2024)to multi-source integration methods (Zhao et al., 2024). These calibration studies predominantly rely on standard question-answering datasets, and it is likely that the LLMs have been trained on a large fraction of the questions used (Tian et al., 2023; Wei et al., 2024; Zhang et al., 2024; Xiong et al., 2024). Thus, these studies do not address LLMs’ confidence in questions that require them to reason. By contrast, we utilize algorithmically generated questions that guarantee minimal training contamination, allowing us to study confidence in answers that involve reasoning by LLMs. We then ask the LLMs to assess the confidence in their answers, closely following similar protocols in human experiments. We elicit the LLM’s confidence in the correctness of the answer, the facts used, and the reasoning. We set the LLM’s temperature to zero, where possible.
@@
We find that all five LLMs we examine are overconfident (Table 1). The LLMs on average overestimate the probability that their answer is correct between 20% (for GPT o1) and 60% (for GPT 3.5), with the other models falling somewhere in between ($p<0.005$ in all cases). We find that more advanced models such as GPT 4o or GPT o1 have higher accuracy than GPT 3.5 or Llama 3.2, in line with earlier findings (Shahriar et al., 2024; Zhong et al., 2024; Dubey et al., 2024). Interestingly, the confidence judgments across models are very similar, and in no way reflect the large differences in accuracy rates.
@@
Next, we examine how, within each LLM model, accuracy is related to confidence across the different questions. Panels A and B of Figure 1 show the accuracy rates as a function of confidence, for the GPT and Llama models, respectively. The panel shows a strong and positive association between accuracy and confidence. The graphs also show that even when a model is fully confident in its answer, there is still substantial bias: Table 2 shows that the bias varies between 15% and 21% for the top performing models (GPT 4o, o1, and Llama 3.1), and even more for GPT 3.5 and Llama 3.2.
@@
For all models except Llama 3.2, bias increases with lower confidence levels. The estimates in Table 2 quantify this effect: a 10% drop in confidence is associated with a drop in accuracy of around 25% for GPT 4o and GPT o1, and around 13% for Llama 3.1 (see Table 2). This implies that the bias gets larger as the models’ confidence declines.
@@
The strong association between confidence and accuracy also extends each component of the reasoning process in which the LLMs have to engage. For each question, we separately elicit the the LLM’s confidence that the facts used are correct, and that the reasoning applied is correct. These confidence measures are highly correlated with confidence that the overall answer is correct (correlations range from 0.49 to 0.92 across models , see Table D.7 and Table D.7 in the SI). We also calculate the semantic similarity between the facts used and the reasoning applied with the sufficient set of facts, and confidence that the reasoning is correct (see section A.3 in the SI for details). We find significant associations between confidence and similarity for, both, facts and reasoning ($p<0.001$ in all cases, see Tables D.7 and D.7 in the SI). Thus, there is a strong, but biased, relationship between accuracy and confidence in the answers, and this association extends to the necessary components for the reasoning.
We also examined the robustness of these results. In a first robustness test, we checked whether committing the model to a response and then asking about the confidence contributes to overconfidence. We set up analogous prompts where, instead of eliciting a direct yes/no answer, and subsequently a confidence statement, we ask directly "what is the probability that the answer is ‘yes’?" We then derive the implied response and confidence from those statements. We also elicit the same measure for the probability of the correct answer being "no." Table S18 in the SI shows that there are no meaningful improvements in accuracy or calibration in either variants of the prompts or any LLM. In several cases, accuracy decreased and bias increased substantially in the alternative prompts. Thus, asking the LLM to first commit to an answer, and then assess the confidence in the answer does not contribute importantly to our baseline findings. Second, we varied the temperature of the LLM, i.e. the extent to which it picks the most likely response (at temperature zero, our baseline temperature) or allows for randomness in the response (at higher temperatures). As Table S19 in the SI shows, there is no quantitatively important effect of the temperature on accuracy or confidence: bias tends to decrease slightly at higher temperatures, by only in the order of a a few percentage points. [^4] We also increased the temperature to 1.5, a value that would be considered too high for most applications. Indeed, most models were unable to produce complete answers at this temperature. The exeception is GPT 3.5, which still manages to produce answers, and enjoys higher accuracy. However, with a baseline accuracy of 35%, even flipping a coin would increase baseline accuracy. We also examine whether varying the temperature and prompt together produced more accurate and better calibrated results, but find little evidence thereof (see section F.2 in the SI for details).
@@
Finally, we also produced five replications of the results for GPT 3.5 and 4o at temperature 0, spanning over several months. We retrieve nearly identical results with regard to the answer and the confidence. The answers change in less than 2 percent in each iteration. The changes we observe also do not have a clear direction: overall, accuracy only increase by 1.7 percent after five replications (see section F.3 in the SI).
@@
We compare the accuracy, confidence and confidence gradient of LLMs to a human benchmark. In a large-scale online experiment on Prolific, we recruit participants to answer 10 questions of a subset of 2000 (see SI section B). Like the LLMs, the human participants have to pick one of the answers (Yes or No), and indicate the probability with which they believe that their answer is correct. Human participants answer 66% of the questions correctly. This puts them on par in terms of accuracy with GPT 4o or Llama 3.1, but slightly behind GPT o1. Their confidence is significantly lower than that of any of the LLMs: human participants think their accuracy is 70%, thus overestimating it by only 4% on average ($p<0.001$, see Table S2 in the SI). This is far lower overconfidence than any of the LLMs with comparable accuracy. There is a maked difference in shape of the distribution of confidence (see Figure S5 in the SI): compared to LLMs, humans’ confidence appears far more nuanced, and with a mode (nearly 35% of the answers) stating that it is 50:50 whether their answers are correct.
@@
Panel C of Figure 1 shows the relationship between accuracy and confidence in the human sample, compared to the three most accurate LLMs. Two striking features emerge from the graph: first, when humans are 100% sure that their answer is correct, their accuracy rate is significantly lower at 81% ($p<0.001$, see column 1 in Table S11 ). Thus, for questions that humans consider relatively easy and are thus 100% sure of their answer, they are highly overconfident. This level of overconfidence is roughly on par with the same conditional accuracy rates of GPT 4o or Llama 3.1. However, while originating from the same level of bias when 100% sure of their answers, the confidence gradients between LLMs and humans are sharply different. As discussed before, LLMs become more biased as they enter more uncertain territory. In sharp contrast, the human confidence gradient is only about 0.5, indicating that human bias decreases as humans become less certain. As can be seen in the figure, humans eventually end up slightly underconfident: when participants think their accuracy is only 50%, it is actually slightly higher (54%, significantly higher than they expect $p<0.01$).
@@
We interpret this sharp difference in the confidence gradient as a stark manifestation of the Dunning-Kruger effect (Dunning, 2011; Kruger and Dunning, 1999). Our test cases require LLMs to reason to answer them, as they are unlikely to be trained on the answer. This may cause LLMs being "trapped" in their prediction model in a way that is hard for them to overcome: the model-internal mechanisms cannot provide them with a sense of what knowledge is not part of their training data. The LLMs pick the answer that is most consistent with what they were trained on, and form an estimate of their accuracy based on what is in their training data. As LLMs move into less certain territory, this problem gets stronger, and causes LLMs to have much less of a sense of the limits of their knowledge. Human participants may recognize in a question, e.g., that one of the names is a Roman emperor, and realize that she does not know much about Roman history. By contrast, in an LLM, there is no mechanism to tell it that there might be other Roman public figures that it wasn’t trained on. This results in overconfidence far stronger than what is observed in humans though the mechanisms of the Dunning-Kruger effect. Interestingly, while more advanced models like GPT 4o and GPT o1 have higher accuracy rates for test cases where they are certain their answer is correct, they also display a significantly stronger confidence gradient, and, hence a stronger Dunning-Kruger effect.
@@ Discussion
The implications of LLM overconfidence span practical, ethical and technical dimensions across stakeholder groups. Our results provide users with a new benchmark for how users of LLMs should approach reasoning from LLMs, counteracting the widespread overestimation of LLM capabilities (Klingbeil et al., 2024; Holbrook et al., 2024; Choudhury and Chaudhry, 2024). Secondly, our results provide guidance for the future development of LLMs. Architectural choices in training objectives currently prioritize fluency over accuracy, incentivizing fabricated but coherent responses (Yin et al., 2023; Lozić and Štular, 2023). The lack of a built-in uncertainty correction mechanism generates hallucinated responses and overconfidence (Shorinwa et al., 2024). Current evaluation metrics (Wei et al., 2024; Geng et al., 2024) fail to capture calibration quality, necessitating new frameworks for quantifying and gauging overconfidence such as the one we presented here. Developing novel interpretability frameworks (Wen et al., 2024; Roitberg et al., 2022) that pinpoint how and why overconfident and biased predictions emerge can support clearer communication of model uncertainty and limitations to users.
@@
Model a) Accuracy b) LLM confidence c) Bias
rate in answer
@@
  • Notes: $N=10,000$ questions submitted to LLM (see 2 for exact number of observations in each LLM). Accuracy rate is the fraction of correct answers. Confidence is elecited from LLM: "What is the probability that your answer is correct? Bias is defined as confidence - accuracy rate. Standard deviations in parentheses (divide by 100 to obtain approximate SE of means). All accuracy rates and confidence measures are significantly different between models. Average bias in each model is significantly different from zero. ($p<0.001$, see SI Table S9.)
@@
Predicted accuracy 0.42 0.79 0.85 0.81 0.65
if fully confident (0.01) (0.01) (0.01) (0.01) (0.01)
NN 9,795 9,973 9,990 9,750 9,910
@@
  • Notes: Dependent variable is correct answer (= 1). The table presents OLS estimates of the gradient of accuracy with respect to self-reported confidence in answer across models (see Table 1 for definition). Confidence levels are normalized by subtracting 1 from the original confidence scores. Robust standard errors in parentheses. All predicted accuracy rates at full confidence are different from 1 at $p<0.005$. Confidence gradients are all greater than one ($p<0.005$), except for Llama 3.2.
Treatment interacted with: Human Baseline Confidence LLM Confidence
Dependent variable: $\Delta$ Accuracy $\Delta$ Bias $\Delta$ Accuracy $\Delta$ Bias
@@
  • Notes: OLS estimates, displaying the difference-in-difference estimates of LLM exposure by different subgroups. Standard errors in parentheses are adjusted for clustering at the participant and question level. $N=1,161$ subjects answered 10 questions each in a intervention study design (see section C in the SI for details of the design). High Confidence is defined as confidence above 0.9 for LLM and above the median for human baseline. The regressions contain a constant term (not shown), and also control for high confidence indicators, though these results are not displayed in the table. Figure S4 visualizes the overall effects in the different subgroups. All main treatment effects are significantly different from control condition ($p<0.005$ in all cases). All treatments have differential impact on high vs. low confidence subgroups ($p<0.005$ in all cases).
@@
Notes: Panel A shows the relationship between accuracy and confidence for GPT models, and Panel B shows the same relationship for Llama models. Panel C compares the accuracy-confidence relationship between advanced LLM models and human responses on the question subset for the human baseline benchmark (see section B of the SI for details on the design). The size of circles in each panel scales with the number of observations, and all slopes are significant ($p<0.001$ in all cases; see Table 2 for Panel A and B, and Table S11 in Supplementary Information for Panel C). Shaded areas indicate 95% confidence bands.
@@
Notes: Panel A shows the treatment effects on change in accuracy across experimental conditions. Panel B shows the treatment effects on the change in bias across the same conditions. Error bars represent 95% confidence intervals. Both treatment conditions have significant effects on accuracy and bias relative to the control group ($p<0.005$ in all cases). LLM Answer + Confidence has a significantly larger effect on on bias than LLM Answer $(p<0.005)$. For more details, see SI Table S12.